IMPROVED POTENTIAL ENERGY CURVE AND MOLECULAR CONSTANTS FOR THE A^{3}Π_{1u} STATE OF I_{2}

\underline{MARJAN MOHAMMADI} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Some years ago we reported an analysis of a high resolution absorption spectrum for the A^{3}Π_{1u} − X^{1}Σ_{g}^{+} system of I_{2} consisting of approximately 9700 lines from 79 bands spanning the vibrational range v′ = 0 − 35 and v′′ = 3 − 17.^{a} Results were presented in the forms both of empirical Dunham-type expansions and of near-dissociation expansion (NDE) expressions for the band constants of the A–state. Because of the use of NDE’s, the resulting expressions were expected to be fairly reliable for extrapolation to dissociation (v_{D} = 55.57(±0.01)). Shortly thereafter, Yukiya et al.^{b} reported additional high resolution absorption data spanning the range v′(A)=16 − 45, and showed that the Appadoo et al.^{a} NDE expression for B_{v} did not, in fact, extrapolate reliably. However, they analyzed their results only in terms of a set of band constants for levels v′(A) = 16 − 45. The present paper describes a comprehensive analysis of both data sets (which together span 99.2% of the potential well) which critically compares the utility of Dunham expansions vs. NDE expressions vs. direct-potential-fit techniques, and yields an optimum description of this system.

a D.R.T. Appadoo, R.J. Le Roy, P.F. Bernath, S. Gerstenkorn, P. Luc, J. Vergès J. Sinzelle, J. Chevillard and Y. D’Aignaux, J.

Chem. Phys. 104, 903–913 (1996).

b T. Yukiya, N. Nishimiya and M. Suzuki, J. Mol. Spectrosc. 182, 271–279 (1997).